OpenStudy (xxlyh):

Will Medal!! What is the solution to the following equation? 2(3x –7) + 18 = 10 1 3 5 6

OpenStudy (xxlyh):

I think it's 1 can someone correct me if im wrong?

OpenStudy (scarlettfarra2000):

Simplifying 2(3x + -7) + 18 = 10 Reorder the terms: 2(-7 + 3x) + 18 = 10 (-7 * 2 + 3x * 2) + 18 = 10 (-14 + 6x) + 18 = 10 Reorder the terms: -14 + 18 + 6x = 10 Combine like terms: -14 + 18 = 4 4 + 6x = 10 Solving 4 + 6x = 10 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4' to each side of the equation. 4 + -4 + 6x = 10 + -4 Combine like terms: 4 + -4 = 0 0 + 6x = 10 + -4 6x = 10 + -4 Combine like terms: 10 + -4 = 6 6x = 6 Divide each side by '6'. x = 1 Simplifying x = 1 You were correct @xxlyh

OpenStudy (xxlyh):

oh thank you!! do you mind helping me correct one more? x)

OpenStudy (scarlettfarra2000):

Sure

OpenStudy (xxlyh):

The steps below show the incomplete solution to find the value of x for the equation 6x + 10 –4x = –4 –11: Step 1: 6x + 10 –4x = –4 –11 Step 2: 6x + 10 –4x = –15 Step 3: 2x + 10 = –15 Which of these is most likely the next step? 2x = 5 2x = –25 2x = 25 2x = 150

OpenStudy (xxlyh):

2x = –25 was my answer

OpenStudy (anthonyym):

The next step in getting x by itself is to subtract 10 from both sides.

OpenStudy (anthonyym):

Yes that's correct.

OpenStudy (xxlyh):

thank you so much for your help!!

OpenStudy (scarlettfarra2000):

Yeah @anthonyym right your answers right sorry double guessing myself